OpenStudy (annjewell1998):

Please help! What is the range of y = 5x -2 if the domain is {-3, -1, 0, 1, 3}? a {-13, -3, 2, 3, 1} b {12, 3, 0, 3, 13} c {-17, -7, -2, 3, 13}

Nnesha (nnesha):

Domain = x values range = y values substitute x for each domain value to get the range

OpenStudy (annjewell1998):

so i'd start with y=5(-3)-2 right?

Nnesha (nnesha):

good!

OpenStudy (annjewell1998):

So y would equal -17

Nnesha (nnesha):

correct now substitute x for the next number in domain just like u did for -3

OpenStudy (annjewell1998):

(-17)=5(-3)-2 (-7)=5(-1)-2 (-2)=5(0)-2 (3)=5(1)-2 (13)=5(3)-2 So the answer is { -17, -7, -2, 3, 13}

Nnesha (nnesha):

yes correct! good job!

OpenStudy (annjewell1998):

Thank you!

Nnesha (nnesha):

my pleasure!